# Federated Learning With Energy Harvesting Devices: An MDP Framework

### v1

@@ VIII-B Proof of Theorem 2
$J^{*}_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})=\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{i=t}^{T}A^{T-i}c(\boldsymbol{h}_{i},\boldsymbol{p}_{i})|\boldsymbol{h}_{t},\boldsymbol{b}_{t}\right],\forall t.$ (64)
Then, we prove by induction that $J^{*}_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})$ is equal to the value function $V_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})$. It is obvious that $J^{*}_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})=V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})$ according to the definition of $V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})$ as shown in (27). Now, suppose that for some $t$, we have $J^{*}_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})=V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})$. Then, we have
$J^{*}_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})=$ $\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{i=t}^{T}A^{T-i}c(\boldsymbol{h}_{i},\boldsymbol{p}_{i})|\boldsymbol{h}_{t},\boldsymbol{b}_{t}\right]$
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\Bigg{\{}A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$
$+\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{i=t+1}^{T}A^{T-i}c(\boldsymbol{h}_{i},\boldsymbol{p}_{i})|\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1}\right]\Bigg{\}}$
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\left\{A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+J^{*}_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})\right\}$
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\left\{A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})\right\}$
@@
$=$ $\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})|\boldsymbol{h}_{1},\boldsymbol{b}_{1}\right].$
Now, we have proved that $V_{1}(\boldsymbol{h}_{1},\boldsymbol{b}_{1})$ is equal to the optimal objective value of Problem 1.
@@ VIII-C Proof of Lemma 1
We first prove that the final-time value function $V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})$ is convex in its battery state $\boldsymbol{b}_{T}$. According to [37, Theorem 9], the packet error rate $q_{n,t}(h_{n,t},p_{n,t})$ is decreasing and convex in $p_{n,t}$. Then, we can obtain that the cost $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$ is decreasing and convex in $\boldsymbol{p}_{t}$. Additionally, at the final time slot $T$, all the remaining energy in the battery should be used up for transmission, i.e., $b_{n,T}=e_{n}^{\text{cmp}}+e_{n,T}^{\text{com}}=e_{n}^{\text{cmp}}+\frac{p_{n,T}s_{n}}{r_{n,T}}$. Thus, we can obtain that the final-time value function $V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})=\min_{\boldsymbol{p}_{T}}\left\{c(\boldsymbol{h}_{T},\boldsymbol{p}_{T})\right\}$ is convex in $\boldsymbol{b}_{T}$. Now assume that $V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1}),1\leq t\leq T-1$, is convex in $\boldsymbol{b}_{t+1}$ for given $\boldsymbol{h}_{t+1}$. Then, we have that function
@@ VIII-D Proof of Theorem 4
According to [29, Theorem 4.7.5], to show that $p_{n,t}^{*}$ is non-increasing in $h_{n,t}$, it suffices to verify that the following four conditions hold:
@@
(3) $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$ is submodular in $h_{n,t}$ and $p_{n,t}$.
@@
First, according to the definition of $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$ in (23), it is evident that $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$ is decreasing in $h_{n,t}$, since $q_{n,t}(h_{n,t},p_{n,t})$ is decreasing in $h_{n,t}$. Thus, condition (1) holds. Moreover, we have $\frac{\partial^{2}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})}{\partial h_{n,t}\partial p_{n,t}}\leq 0$, which indicates that $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$ is a submodular function in $h_{n,t}$ and $p_{n,t}$, and thus condition (3) holds.
@@ VIII-E Proof of Theorem 5
Recall that $\overline{\Omega}_{t}$ denotes the set of scheduled devices under the optimal relaxed policy $\overline{\pi}^{*}$ at time slot $t$. According to the truncation procedure, if $|\overline{\Omega}_{t}|>R$, $R$ devices will be chosen randomly from the set $\overline{\Omega}_{t}$ and scheduled. Therefore, the probability that a device $n$ belongs to the set $\overline{\Omega}_{t}$ but is not scheduled can be computed as $\frac{(|\overline{\Omega}_{t}|-R)^{+}}{|\overline{\Omega}_{t}|}$, where $(\cdot)^{+}=\max\left\{\cdot,0\right\}$. At time slot $t$, the per-device cost of device $n$ is no more than $\frac{\eta\rho}{2ND_{n}}$. Then, the gap between the cumulative costs under truncation policy $\tilde{\pi}$ and the optimal relaxed policy $\overline{\pi}^{*}$ is upper bounded by
$J(\tilde{\pi})-J(\overline{\pi}^{*})\leq$ $\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[\sum_{t=1}^{T}\sum_{n=1}^{N}\frac{\eta\rho(|\overline{\Omega}_{t}|-R)^{+}}{2ND_{n}|\overline{\Omega}_{t}|}\right]$ (69)
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(a)}}}}{{\leq}}}$ $\frac{\eta\rho}{2D_{\textup{min}}}\sum_{t=1}^{T}\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[\frac{(|\overline{\Omega}_{t}|-R)^{+}}{|\overline{\Omega}_{t}|}\right]$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(b)}}}}{{\leq}}}$ $\frac{\eta\rho}{2RD_{\textup{min}}}\sum_{t=1}^{T}\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[(|\overline{\Omega}_{t}|-R)^{+}\right]$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(c)}}}}{{\leq}}}$ $\frac{\eta\rho}{2RD_{\textup{min}}}\sum_{t=1}^{T}\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[\left(|\overline{\Omega}_{t}|-\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[|\overline{\Omega}_{t}|\right]\right)^{+}\right]$
$\leq$ $\frac{\eta\rho}{2RD_{\textup{min}}}\sum_{t=1}^{T}\underbrace{\mathbb{E}_{\overline{\pi}^{*}\!}\!\left[\left||\overline{\Omega}_{t}|-\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[|\overline{\Omega}_{t}|\right]\right|\right]}_{=\textup{MAD}(|\overline{\Omega}_{t}|)},$
where (a) is due to $D_{\textup{min}}=\min_{n}\left\{D_{n}\right\}$, (b) follows from $\frac{(|\overline{\Omega}_{t}|-R)^{+}}{|\overline{\Omega}_{t}|}\leq\frac{(|\overline{\Omega}_{t}|-R)^{+}}{R}$, and (c) holds as $\mathbb{E}_{\overline{\pi}^{*}}\left[|\overline{\Omega}_{t}|\right]\leq R$. The notation $\textup{MAD}({X}):=\mathbb{E}[|X-\mathbb{E}[X]|]$ stands for the mean absolute deviation of its argument.
Recall that $\overline{\Omega}_{t}=\left\{n\,|\,\overline{\beta}_{n,t}=1,n\in[N]\right\}$ and thus $|\overline{\Omega}_{t}|=\sum_{n=1}^{N}\overline{\beta}_{n,t}$. The scheduling indicators $\left\{\overline{\beta}_{n,t},n\in[N]\right\}$ under the optimal relaxed policy $\overline{\pi}^{*}$ are independent binary random variables, since the optimal relaxed policy is obtained by decoupling the relaxed problem into per-device subproblems. Thus, we have that the standard deviation of $|\overline{\Omega}_{t}|$, $\textup{STD}(|\overline{\Omega}_{t}|)$, is upper bounded by
@@
According to [38], we have that $\textup{MAD}(X)\leq\textup{STD}(X)$ for arbitrary random variable $X$. Then, combining (70) into (69), we have
$J(\tilde{\pi})-J(\overline{\pi}^{*})\leq\frac{\eta\rho T\sqrt{N}}{2RD_{\textup{min}}}=\frac{\eta\rho T}{2\kappa D_{\textup{min}}\sqrt{N}}.$ (71)
@@
$J(\tilde{\pi})-J(\pi^{*})\leq$ $J(\tilde{\pi})-J(\overline{\pi}^{*})\leq\frac{\eta\rho T}{2\kappa D_{\textup{min}}\sqrt{N}}.$ (72)

### v2

@@ VIII-B Proof of Theorem 2
$J^{*}_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})=\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{i=t}^{T}c(\boldsymbol{h}_{i},\boldsymbol{p}_{i})|\boldsymbol{h}_{t},\boldsymbol{b}_{t}\right],\forall t.$ (67)
Then, we prove by induction that $J^{*}_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})$ is equivalent to the value function $V_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})$. It is obvious that $J^{*}_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})=V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})$ according to the definition of $V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})$ as shown in (29). Now, suppose that for some $t$, we have $J^{*}_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})=V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})$. Then, we have
$J^{*}_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})$
$=$ $\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{i=t}^{T}c(\boldsymbol{h}_{i},\boldsymbol{p}_{i})|\boldsymbol{h}_{t},\boldsymbol{b}_{t}\right]$
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\Bigg{\{}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{i=t+1}^{T}c(\boldsymbol{h}_{i},\boldsymbol{p}_{i})|\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1}\right]\Bigg{\}}$
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\left\{c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+J^{*}_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})\right\}$
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\left\{c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})\right\}$
@@
$=$ $\min_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})|\boldsymbol{h}_{1},\boldsymbol{b}_{1}\right].$
Now, we have proved that $V_{1}(\boldsymbol{h}_{1},\boldsymbol{b}_{1})$ is equivalent to the optimal objective value of Problem 1.
@@ VIII-C Proof of Lemma 1
We first prove that the final-time value function $V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})$ is convex in its battery state $\boldsymbol{b}_{T}$. According to [39, Theorem 9], the packet error rate $q_{n,t}(h_{n,t},p_{n,t})$ is decreasing and convex in $p_{n,t}$. Then, we have that the cost $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$ is decreasing and convex in $\boldsymbol{p}_{t}$. Additionally, at the final time slot $T$, all the remaining energy in the battery should be used up for transmission, i.e., $b_{n,T}=e_{n}^{\text{cmp}}+\frac{p_{n,T}s_{n}}{r_{n,T}}$. Thus, we can obtain that the final-time value function $V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})=\min_{\boldsymbol{p}_{T}}\left\{c(\boldsymbol{h}_{T},\boldsymbol{p}_{T})\right\}$ is convex in $\boldsymbol{b}_{T}$. Now assume that $V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1}),1\leq t\leq T-1$, is convex in $\boldsymbol{b}_{t+1}$ for given $\boldsymbol{h}_{t+1}$. Then, we have that function
@@ VIII-D Proof of Theorem 4
For clarity, given an action $\boldsymbol{p}_{t}$, we define $\boldsymbol{p}_{n,t}^{+}=\{p_{1,t},\ldots,p_{n,t}+1,\ldots,p_{N,t}\}$. According to [40, Theorem 1], to show that $p_{n,t}^{*}$ is non-increasing in $h_{n,t}$, it suffices to verify that the following four conditions hold:
@@
(3) There exists a function $\gamma_{\boldsymbol{h}_{t}^{-},\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{t}}>0$ and non-decreasing in $\boldsymbol{p}_{t}$ such that for any $\boldsymbol{p}_{t}$ and $\boldsymbol{h}_{t}^{-}\leq\boldsymbol{h}_{t}^{+}$, we have $c(\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{n,t}^{+})-c(\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{t})\leq\gamma_{\boldsymbol{h}_{t}^{-},\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{t}}[c(\boldsymbol{h}_{t}^{-},\boldsymbol{p}_{n,t}^{+})-c(\boldsymbol{h}_{t}^{-},\boldsymbol{p}_{t})]$.
@@
Note that condition (3) is less restrictive than the submodularity condition (33), and when $\gamma_{\boldsymbol{h}_{t}^{-},\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{t}}=1$, condition (3) is equivalent to the submodularity condition.
First, it is clear that condition (1) holds due to the fact that the packet error rate $q_{n,t}(h_{n,t},p_{n,t})$ is decreasing in $h_{n,t}$. Moreover, let $\gamma_{\boldsymbol{h}_{t}^{-},\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{t}}=\frac{c(\boldsymbol{h}_{t}^{+},\boldsymbol{p}_{t})-c(\boldsymbol{h}_{t}^{+},\boldsymbol{p}^{+}_{n,t})}{c(\boldsymbol{h}_{t}^{-},\boldsymbol{p}_{t})}>0$, and then we can obtain that condition (3) holds.
@@ VIII-E Proof of Theorem 5
Recall that $\overline{\Omega}_{t}$ denotes the set of scheduled devices under the optimal relaxed policy $\overline{\pi}^{*}$. According to the truncation procedure, if $|\overline{\Omega}_{t}|>R$, $R$ devices will be chosen randomly from the set $\overline{\Omega}_{t}$ and scheduled. Therefore, the probability that a device $n$ belongs to the set $\overline{\Omega}_{t}$ but is not scheduled is $\frac{(|\overline{\Omega}_{t}|-R)^{+}}{|\overline{\Omega}_{t}|}$, where $(\cdot)^{+}=\max\left\{\cdot,0\right\}$. At time slot $t$, the per-device cost of device $n$ is no more than $\frac{2LG^{2}}{N\sqrt{KT}}$. Then, the gap between the cumulative costs under truncation policy $\tilde{\pi}$ and the optimal relaxed policy $\overline{\pi}^{*}$ is upper bounded by
$J(\tilde{\pi})-J(\overline{\pi}^{*})\leq$ $\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[\sum_{t=1}^{T}\sum_{n=1}^{N}\frac{2LG^{2}(|\overline{\Omega}_{t}|-R)^{+}}{N\sqrt{KT}|\overline{\Omega}_{t}|}\right]$
$=$ $\frac{2LG^{2}}{\sqrt{KT}}\sum_{t=1}^{T}\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[\frac{(|\overline{\Omega}_{t}|-R)^{+}}{|\overline{\Omega}_{t}|}\right]$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(a)}}}}{{\leq}}}$ $\frac{2LG^{2}}{R\sqrt{KT}}\sum_{t=1}^{T}\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[(|\overline{\Omega}_{t}|-R)^{+}\right]$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(b)}}}}{{\leq}}}$ $\frac{2LG^{2}}{R\sqrt{KT}}\sum_{t=1}^{T}\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[\left(|\overline{\Omega}_{t}|-\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[|\overline{\Omega}_{t}|\right]\right)^{+}\right]$
$\leq$ $\frac{2LG^{2}}{R\sqrt{KT}}\sum_{t=1}^{T}\underbrace{\mathbb{E}_{\overline{\pi}^{*}\!}\!\left[\left||\overline{\Omega}_{t}|-\mathbb{E}_{\overline{\pi}^{*}\!\!}\left[|\overline{\Omega}_{t}|\right]\right|\right]}_{=\textup{MAD}(|\overline{\Omega}_{t}|)},$ (72)
where (a) follows from $\frac{(|\overline{\Omega}_{t}|-R)^{+}}{|\overline{\Omega}_{t}|}\leq\frac{(|\overline{\Omega}_{t}|-R)^{+}}{R}$, and (b) holds as $\mathbb{E}_{\overline{\pi}^{*}}\left[|\overline{\Omega}_{t}|\right]\leq R$. The notation $\textup{MAD}({X}):=\mathbb{E}[|X-\mathbb{E}[X]|]$ stands for the mean absolute deviation of its argument.
Let $\overline{\beta}_{n,t}$ denote the scheduling indicator under the optimal relaxed policy $\overline{\pi}^{*}$, and we have $|\overline{\Omega}_{t}|=\sum_{n=1}^{N}\overline{\beta}_{n,t}$. The scheduling indicators $\left\{\overline{\beta}_{n,t},n\in[N]\right\}$ are independent binary random variables, since the optimal relaxed policy is obtained by decoupling the relaxed problem into per-device subproblems. Thus, we have that the standard deviation of $|\overline{\Omega}_{t}|$, $\textup{STD}(|\overline{\Omega}_{t}|)$, is upper bounded by
@@
According to [41], we have that $\textup{MAD}(X)\leq\textup{STD}(X)$ for arbitrary random variable $X$. Then, combining (73) into (VIII-E), we have
$J(\tilde{\pi})-J(\overline{\pi}^{*})\leq\frac{2LG^{2}\sqrt{NT}}{R\sqrt{K}}=\frac{2LG^{2}\sqrt{T}}{\kappa\sqrt{NK}}.$ (74)
@@
$J(\tilde{\pi})-J(\pi^{*})\leq$ $J(\tilde{\pi})-J(\overline{\pi}^{*})\leq\frac{2LG^{2}\sqrt{T}}{\kappa\sqrt{NK}}.$ (75)